# Accretion onto supermassive and intermediate mass black holes in cosmological simulations

### v1

@@ 3.2 Intermediate mass black hole accretion at high redshift
We now use the models that best match the global BHMD in our large volume simulation and apply them with the same parameters to study their behaviour at high redshift and for intermediate mass black holes. We analyse the high redshift, smaller volume L12.5n768 simulations in which the only other modifications are the seeding parameters: the seeding halo mass is lowered by a factor of to M⊙ and the seed black hole mass by a factor to M⊙. Lowering the seed halo mass less than the seed mass is motivated by avoiding over-seeding compared to physically motivated seed models (see Bhowmick et al. 2024b, for studies using the IllustrisTNG model). We note that our main aim in this work is not to study realistic populations of high redshift IMBHs, but only to study the relative differences that different accretion models have on the high redshift IMBH population.
@@
Figure 6 shows the BHARD and BHMD as a function of redshift for the IMBH population. In contrast to the SMBH population shown in Figure 2, the BHARD differs by orders of magnitude depending on the accretion model. This is a direct consequence of the different scaling with black hole mass in the different accretion rate estimates. Notably, however, the BHMD is very similar, only gradually diverging starting at . This is due to the mass growth in some of these models being dominated by seeding, posing a lower limit. Only the ff model shows growth via gas accretion at similar rates. We note, however, that the frequency of seeding in our model is not well constrained and in its redshift dependence likely incorrect (see Tremmel et al. 2017, for a comparison between a gas-based and halo based seeding frequency).
@@
of IMBH mergers until for the 3 models (since the accretion disks make practically no difference in the mass functions, we do not show them here). It should be noted that, due to the chosen seeding parameters, the number of mergers is relatively limited (of order 30 per simulation). Furthermore, we operate with an instantaneous merger model, not taking dynamical friction on spatially resolved scales nor the complex inspiral process on unresolved scales and its delays into account (see Kelley et al. 2017, for a more thorough modelling). Thus, the result should not be interpreted as an absolute prediction, but only the relative changes due to accretion model variation are relevant. Indeed, the models with a steep mass function tng and mod ff have a merger distribution strongly peaked at the seed mass, with few mergers of higher mass black holes, while the ff model, with its considerable mass growth via accretion has a broader chirp mass distribution.
@@ 4 Discussion
We study the effects of different accretion models in cosmological simulations for both SMBH and IMBH gas accretion. While the models are inspired by high-resolution simulations and analytic considerations, they are intentionally built to be able to parameterize different scalings and represent different mechanisms driving the gas inflow. For the present work, we adjust the parameters to roughly match the SMBH mass density at low redshift (Fig. 2). We analyze the remaining differences in SMBH growth over cosmic time and explore how these parameters would affect the growth of IMBHs at high redshift.
The prefactor in the ff model, is somewhat smaller than in other works using similar models (Ehlert et al. 2023; Wellons et al. 2023), however, we note our larger accretion region compared to cosmological zoom simulations and lack of other criteria such as only accreting sufficiently cold gas, might explain the need for a lower efficiency. In the mod ff model, should be interpreted as the radius dependent mass inflow rate slope not reaching all the way to the Schwarzschild radius, but transitioning to a constant inflow rate with radius at . Note that since , the scaling of the inflow rate with black hole mass changes as , driving the differences between mod ff () and ff (). The Bondi accretion rate scales as and is additionally modified by a sound speed dependence and a different way the density is estimated, making the direct comparison more difficult.
@@ 4.1 Non-uniqueness of the MBH gas accretion in a galaxy formation context
In the SMBH regime, the quite substantial changes of the accretion model have only moderate impacts on the MBHs over cosmic time once the normalization has been set to roughly match the black hole mass density at . While the qualitative shape of the BHARD is largely dictated by the gas accretion rate onto the galaxy (van de Voort et al. 2011), the precise shape changes with mass scaling (Fig. 2), in particular the redshift of peak AGN activity and the magnitude of the reduction to .
@@
The black hole mass functions at are relatively similar. At and , differences are present at the high mass end (Fig. 3, left column). The dominance of mergers for the mass growth of high mass SMBHs at low redshift (Weinberger et al. 2018) likely erases some of these differences by . This is, of course, to a certain degree a consequence of the calibration of the model against the BHMD; i.e. the integrated mass function. Consequently, the black hole mass - stellar mass and black hole mass - velocity dispersion relations (Fig. 5) differ only mildly with slightly flatter relations for the ff model. Yet, given large observational uncertainties and biases, it is hard to imagine how they would be suited to reveal the nature of gas inflow. In this context, it is important to keep in mind that we keep all other parameters constant. A shallower mass dependence, when normalized to a reasonable BHMD, does not produce overmassive black holes relative to the relations at any redshift. Thus, with the model variations shown here, we are unable to reproduce the scaling inferred by Pacucci et al. (2023), nor provide a sufficient scatter to explain the reported detections (Li et al. 2024). Redshift dependent scalings e.g. in the radiative efficiency or a more efficient quasar mode feedback that more efficiently regulates accretion after some MBHs have grown overmassive might be needed to simultaneously produce overmassive high redshift black holes at while not overgrowing them toward .
@@ 4.2 IMBH gas accretion
Unlike in the SMBH regime to which the parameters were calibrated, the different models cause massively different outcomes assuming M⊙ seeds. The resulting BHARD changes by orders of magnitude for the different models (Fig. 6). The resulting bolometric luminosity and Eddington ratio distributions are orders of magnitude apart (Fig. 7), leading also to substantially different occupation fractions provided the threshold luminosity is low enough (Fig. 8). The employed X-ray luminosity thresholds are lower than expected detection limits with Athena and AXIS. Future surveys with these telescopes can detect AGNs with soft ( keV) X-ray luminosities down to erg s-1 and erg s-1 at , respectively (Marchesi et al. 2020). If we used thresholds more in line with (future) X-ray detectable AGNs, the occupation fraction in our boxes would be vanishing for all models. However, we note that this is merely an artifact of the small volume, and consequently low-mass host galaxies simulated here. Simulating larger volumes would naturally produce more massive host galaxies as well, with slightly more massive IMBHs (of mass M⊙ at ). Since the Eddington ratio scales differently with black hole mass for the different models ( for ff, for mod ff and for tng), the differences in Eddington ratio distribution would be less pronounced for more massive IMBHs (assuming no other effects). However, the differences might still be present, producing a sweet-spot between detectability and differences in accretion rate in this mass range. We thus argue that future studies of the detectability of AGNs powered by IMBHs, not just in X-ray but also in optical wavelengths (e.g. Cann et al. 2019) need to carefully take uncertainties in the accretion rate estimate into account.
@@
Interestingly, the dramatic change in accretion rate does not in all cases carry over to the black hole mass density, indicating that its buildup has a floor set by seed formation. For the more optimistic accretion scenarios, however, the mass growth via gas accretion of individual black holes can be substantial, flattening the black hole mass function (Fig. 7, left column). This accretion driven growth also modifies the frequency of IMBH mergers of given chirp mass (Fig. 9), highlighting the importance to take it into consideration when interpreting merger rates, e.g., from future LISA gravitational wave detections (Sesana et al. 2007; Amaro-Seoane et al. 2017). On top of this, uncertainties in the binary inspiral phases (Kelley et al. 2017; Siwek et al. 2020) need to be considered, which we neglect in this work due to the instant merger assumption for MBH binaries.
@@
It is important to note here that the presented seeding is just a simple change in the parameters of the IllustrisTNG seeding prescription and not a prediction for any specific pathway (see Volonteri 2010, for a review). However, it provides a reasonable environment to study the potential for growth of IMBHs via accretion. Considering the BHMD in our simulation, accretion only represents a fraction of the seeded black holes, and a more carefully considered IMBH seeding prescription is needed to fully assess the fractional growth via accretion on these systems along with previously discussed observational constraints. Improved MBH formation models should accurately predict seed locations, redshifts (Bellovary et al. 2011), frequency (Habouzit et al. 2016) and seed mass function (e.g. as modelled in Ni et al. 2022).
@@ 4.3 The role of accretion disks
Apart from varying the accretion model, we also run the different inflow models with an additional model for unresolved accretion disks. Its main function is as a gas reservoir that is filled by the gas inflow and drained over a given timescale. The natural expectation of such a model is that fluctuations in the instantaneous inflow rate will be smoothed out, while the overall accretion rate should not be affected since it is limited by the inflow rate, not the accretion disk (with the exception of the Eddington-limited regime). Indeed for the mass growth, we see very little difference between the mass growth with and without an accretion disk model, only some initial delay due to the accretion disk timescale at the highest redshifts. In quantities proportional to the instantaneous accretion rate, i.e. luminosity and Eddington ratio, no substantial modifications arise in the SMBH regime. In the high redshift IMBH regimes, larger differences are visible. We speculate that this is more related to the higher redshifts and correspondingly shorter structure formation timescales (becoming comparable to the accretion disk timescale) rather than to the mass of the MBHs.
@@ 5 Summary and conclusions
We present a new, simple model for massive black hole accretion and compare its effect to the IllustrisTNG model. The key aspects of this model are:
@@
The inflow estimate is based on a free-fall time which scales more weakly with black hole mass than the Bondi formula and linearly with enclosed gas mass.
@@
The surrounding gas properties are estimated from gas within a fixed proper distance from the black hole.
@@
A model for a radial dependence of the unresolved part of the mass inflow rate is introduced. Its scaling affects the overall black hole mass scaling.
@@
The models’ normalization parameter is chosen to result in a roughly comparable black hole mass density. Having fixed this parameter, the key effects in the SMBH population are:
@@
At fixed final BHMD, the BHARD can vary only moderately. Interestingly, the differences at are very small when keeping seeding and AGN feedback the same. The BHARD at cosmic noon vs. at can vary somewhat depending on accretion rate estimate, with a less pronounced peak and less dimming toward for less efficient accretion (in the free-fall based model in our case).
@@
The quasar luminosity functions are remarkably similar for the different accretion models. The main effect, most obvious in the occupation fractions at , is a non-linear interaction with AGN feedback. The transition to efficient AGN feedback, which differs for the different accretion models, has a more substantial effect on luminous AGN than the inflow rate estimate (within the bounds tested in this work). In future work, a recalibration of the AGN feedback is needed to match galaxy population constraints to properly assess the uncertainties in the MBH population. Notably, even prior to AGN feedback becoming a relevant factor for the galactic star formation rate, our results indicate an AGN feedback induced self-regulation of the accretion rate.
@@
The black hole mass - stellar mass and black hole mass - velocity dispersion relations change moderately, with a slight flattening of both relations for the free-fall based model. Notably, provided the overall normalization is calibrated to the mass density, a weaker mass dependence of the accretion rate does not lead to substantially overmassive black holes in the low mass end of the black hole mass-stellar mass relation, a trend observations suggest at high redshift (Pacucci et al. 2023). We speculate that a variable radiative or feedback efficiency would be required to achieve this while not violating constraints for the X-ray background.
We therefore conclude that the precise accretion rate formula used in galaxy formation simulations has only minor impacts on the galaxy and the SMBH populations once tuned to match the black hole mass density at . However, when we apply the different accretion rate estimates to a high-redshift simulation modeling an IMBH population with seed mass around , drastic changes are evident:
@@
Accretion rate densities differ by 4 orders of magnitude, while the mass growth is less affected due to growth via seeding. This indicates that our lack of understanding of the accretion mechanism leads to substantial uncertainties for possible signatures and growth histories of high redshift IMBHs.
@@
AGN luminosities differ by orders of magnitude depending on the model. We only show occupation fractions with luminosity thresholds below the detection limit of future X-ray surveys with Athena or AXIS (Marchesi et al. 2020). This is due to our relatively small simulation volume of Mpc side length that only contains low-mass IMBHs at this redshift. However, the trend indicates that slightly more massive black holes that would be present in larger volumes would be detectable at depending on accretion rate estimate. More work with simulations of larger volume is needed to explore if these AGNs could constrain the inflow rate scaling onto MBHs, as well as how sensitive these object are to variations of quasar mode feedback.
These changes highlight the need to understand accretion, in particular the black hole mass dependence of gas inflow to accurately model the high redshift IMBH regime. While we kept the seeding model fixed for the different SMBH simulations, future simulations that model the cosmic evolution of MBHs from light seeds will likely also show substantial accretion model dependent growth in SMBHs caused by the modified early growth history. Further studies combining this effort with effective seed models (such as Bhowmick et al. 2024b) are clearly needed to explore the possible IMBH evolution scenarios at high redshift and their consequences for luminous AGN and IMBH merger events, and the MBH population in the local Universe.
Overall, this new model represents a general gas inflow formula that parameterizes different black hole mass dependences and normalizations. It was developed to match the sprit of the “Learning the Universe” project; i.e., to be applied as an effective model with parameters to be constrained by data or marginalized over in order to infer galaxy formation and cosmological parameters from structure formation. We deliberately choose to only use reliable gas properties in the simulation that are weakly dependent on other model choices such as the employed interstellar medium modelling (see Sivasankaran et al. 2022, for issues with the Bondi estimate). Measuring the simulation quantities at a fixed aperture also allows the model to be tested using high-resolution inflow simulations (Guo et al. 2023, 2024).

### v2

@@ 3.2 IMBH accretion at high redshift
In this step of the study, we use the models that best match the global BHMD in our large volume simulation and apply them with the same parameters to study their behaviour at high redshift and for IMBHs. We analysed the high redshift, smaller volume L12.5n768 simulations in which the only other modifications are the seeding parameters: the seeding halo mass is lowered by a factor of to M⊙ and the seed BH mass by a factor to M⊙. The choice to lower the seed halo mass below that of the seed mass is meant to avoid over-seeding, as compared to physically motivated seed models (see Bhowmick et al. 2024b, for studies using the IllustrisTNG model). We note that our main aim in this work is not to study realistic populations of high redshift IMBHs, but only to study the relative differences that different accretion models have on the high redshift IMBH population.
@@
Figure 6 shows the BHARD and BHMD as a function of redshift for the IMBH population. In contrast to the SMBH population shown in Figure 2, the BHARD differs by orders of magnitude depending on the accretion model. This is a direct consequence of the different scaling with BH mass in the different accretion rate estimates. Notably, however, the BHMD is very similar, only gradually diverging starting at . This is due to the mass growth in some of these models being dominated by seeding, posing a lower limit. Only the ff model shows growth via gas accretion at similar rates. We note, however, that the frequency of seeding in our model is not well constrained and in its redshift dependence likely incorrect (see Tremmel et al. 2017, for a comparison between a gas-based and halo based seeding frequency).
@@
of IMBH mergers until for the three models (since the accretion disks make practically no difference in the mass functions, we do not show them here). It should be noted that, due to the chosen seeding parameters, the number of mergers is relatively limited (of order 30 per simulation). Furthermore, we operated with an instantaneous merger model, not taking dynamical friction on spatially resolved scales nor the complex inspiral process on unresolved scales and its delays into account (see Kelley et al. 2017, for a more thorough modelling). Thus, the result should not be interpreted as an absolute prediction, but only the relative changes due to accretion model variation are relevant. Indeed, the models with a steep mass function tng and mod ff have a merger distribution strongly peaked at the seed mass, with few mergers of higher mass BHs, while the ff model, with its considerable mass growth via accretion has a broader chirp mass distribution.
@@ 4 Discussion
Here, we study the effects of different accretion models in cosmological simulations for both SMBH and IMBH gas accretion. While the models are inspired by high-resolution simulations and analytic considerations, they are intentionally built to be able to parametrise different scalings and represent different mechanisms driving the gas inflow. For the present work, we adjust the parameters to roughly match the SMBH mass density at low redshift (Fig. 2). We analyse the remaining differences in SMBH growth over cosmic time and explore how these parameters would affect the growth of IMBHs at high redshift.
The pre-factor in the ff model, is somewhat smaller than in other works using similar models (Ehlert et al. 2023; Wellons et al. 2023), however, we note our larger accretion region compared to cosmological zoom simulations and lack of other criteria such as only accreting sufficiently cold gas, might explain the need for a lower efficiency. In the mod ff model, should be interpreted as the radius dependent mass inflow rate slope not reaching all the way to the Schwarzschild radius, but transitioning to a constant inflow rate with radius at . Note that since , the scaling of the inflow rate with BH mass changes as , driving the differences between mod ff () and ff (). The Bondi accretion rate scales as and is additionally modified by a sound speed dependence and a different way the density is estimated, making the direct comparison more difficult.
@@ 4.1 Non-uniqueness of the MBH gas accretion in a galaxy formation context
In the SMBH regime, the substantial changes made to the accretion model have only moderate impacts on the MBHs over cosmic time once the normalisation has been set to roughly match the BHMD at . While the qualitative shape of the BHARD is largely dictated by the gas accretion rate onto the galaxy (van de Voort et al. 2011), the precise shape changes with mass scaling (Fig. 2), in particular the redshift of peak AGN activity and the magnitude of the reduction to .
@@
The BH mass functions at are relatively similar. At and , differences are present at the high mass end (Fig. 3, left column). The dominance of mergers for the mass growth of high mass SMBHs at low redshift (Weinberger et al. 2018) likely erases some of these differences by . This is, of course, to a certain degree a consequence of the calibration of the model against the BHMD; namely, the integrated mass function. Consequently, the BH mass-stellar mass and BH mass-velocity dispersion relations (Fig. 5) differ only mildly with slightly flatter relations for the ff model. Yet, given large observational uncertainties and biases, it is hard to imagine how they would be suited to reveal the nature of gas inflow. In this context, it is important to keep in mind that we keep all other parameters constant. A shallower mass dependence, when normalised to a reasonable BHMD, does not produce over-massive BHs relative to the relations at any redshift. Thus, with the model variations shown here, we are unable to reproduce the scaling inferred by Pacucci et al. (2023), nor provide a sufficient scatter to explain the reported detections (Li et al. 2025). Redshift dependent scalings e.g. in the radiative efficiency or a more efficient quasar mode feedback that more efficiently regulates accretion after some MBHs have grown overmassive might be needed to simultaneously produce over-massive high-redshift BHs at while not overgrowing them toward .
@@ 4.2 IMBH gas accretion
Unlike in the SMBH regime to which the parameters were calibrated, the different models cause massively different outcomes assuming M⊙ seeds. The resulting BHARD changes by orders of magnitude for the different models (Fig. 6). The resulting bolometric luminosity and Eddington ratio distributions are orders of magnitude apart (Fig. 7), leading also to substantially different occupation fractions provided the threshold luminosity is low enough (Fig. 8). The employed X-ray luminosity thresholds are lower than expected detection limits with Athena and AXIS. Future surveys with these telescopes can detect AGNs with soft ( keV) X-ray luminosities down to erg s-1 and erg s-1 at , respectively (Marchesi et al. 2020). If we used thresholds more in line with (future) X-ray detectable AGNs, the occupation fraction in our boxes would be vanishing for all models. However, we note that this is merely an artifact of the small volume, and consequently low-mass host galaxies simulated here. Simulating larger volumes would naturally produce more massive host galaxies as well, with slightly more massive IMBHs (of mass M⊙ at ). Since the Eddington ratio scales differently with BH mass for the different models ( for ff, for mod ff and for tng), the differences in Eddington ratio distribution would be less pronounced for more massive IMBHs (assuming no other effects). However, the differences might still be present, producing a sweet-spot between detectability and differences in accretion rate in this mass range. We thus argue that future studies of the detectability of AGNs powered by IMBHs, not just in X-ray but also in optical wavelengths (e.g. Cann et al. 2019) need to carefully take uncertainties in the accretion rate estimate into account.
We note that our simulations are not designed to model gas inflows onto individual IMBHs in great detail, but rather to explore the consequences of different assumptions on a larger population of IMBHs. Dependency on the IMBH formula, however, highlights the importance of dedicated studies of high redshift gas accretion onto IMBH. A dependence on inflow model implies that resolved scales do not include the bottleneck of gas accretion, and smaller scales play a crucial role. With sufficient resolution to capture these scales, however, it is expected that the accretion rate converges independently of accretion formula. Gordon et al. (2024), for example, find a converged accretion independent of accretion scheme for M⊙ IMBHs with simulations reaching resolutions of around pc, i.e., at significantly higher resolution than achievable with uniform volume approaches.
@@
Interestingly, the dramatic change in accretion rate does not in all cases carry over to the BHMD, indicating that its buildup has a floor set by seed formation. For the more optimistic accretion scenarios, however, the mass growth via gas accretion of individual BHs can be substantial, flattening the BH mass function (Fig. 7, left column). This accretion driven growth also modifies the frequency of IMBH mergers of given chirp mass (Fig. 9), highlighting the importance to take it into consideration when interpreting merger rates, for instance, from future LISA gravitational wave detections (Sesana et al. 2007; Amaro-Seoane et al. 2017). In addition, uncertainties in the binary inspiral phases (Kelley et al. 2017; Siwek et al. 2020) need to be considered, which we neglected in this work due to the instant merger assumption for MBH binaries.
However, for reliable predictions of the growth history, dynamical friction and BH recoil kicks are two aspects that are missing in the employed model, which are also of particular importance for IMBHs. Since dynamical friction is relatively inefficient at low masses, IMBHs are not necessarily residing in the centre of galaxies but rather in lower density outskirts, leading to lower accretion rates and consequently mass growth. Also, growth via mergers can be impeded via recoil kicks, which can lead to merged IMBHs to escape the host halo. We leave the study of these effects to future work.
@@
It is important to note here that the presented seeding is just a simple change in the parameters of the IllustrisTNG seeding prescription and not a prediction for any specific pathway (see Volonteri 2010, for a review). However, it provides a reasonable environment to study the potential for growth of IMBHs via accretion. Considering the BHMD in our simulation, accretion only represents a fraction of the seeded BHs, and a more carefully considered IMBH seeding prescription is needed to fully assess the fractional growth via accretion on these systems along with previously discussed observational constraints. Improved MBH formation models should accurately predict seed locations, redshifts (Bellovary et al. 2011), frequency (Habouzit et al. 2016) and seed mass function (e.g. as modelled in Ni et al. 2022).
@@ 4.3 The role of accretion disks
Apart from varying the accretion model, we also ran the different inflow models with an additional model for unresolved accretion disks. Its main function is as a gas reservoir that is filled by the gas inflow and drained over a given timescale. The natural expectation of such a model is that fluctuations in the instantaneous inflow rate will be smoothed out, while the overall accretion rate should not be affected since it is limited by the inflow rate, not the accretion disk (with the exception of the Eddington-limited regime). Indeed for the mass growth, we see very little difference between the mass growth with and without an accretion disk model, only some initial delay due to the accretion disk timescale at the highest redshifts. In quantities proportional to the instantaneous accretion rate, i.e. luminosity and Eddington ratio, no substantial modifications arise in the SMBH regime. In the high redshift IMBH regimes, larger differences are visible. We speculate that this is more related to the higher redshifts and correspondingly shorter structure formation timescales (becoming comparable to the accretion disk timescale) rather than to the mass of the MBHs.
@@ 5 Summary and conclusions
We present a new, simple model for MBH accretion. We compare its performance and results to those of the IllustrisTNG model. The key aspects of our model are:
@@
The inflow estimate is based on a free-fall time, which scales more weakly with BH mass than the Bondi formula. It also scales linearly with an enclosed gas mass.
@@
The surrounding gas properties are estimated from gas within a fixed proper distance from the BH.
@@
A model for a radial dependence of the unresolved part of the mass inflow rate is also introduced. Its scaling affects the overall BH mass scaling.
@@
The normalisation parameter in these models was chosen to result in a roughly comparable BHMD. Fixing this parameter, the key effects in the SMBH population are:
@@
At fixed final BHMD, the BHARD can vary only moderately. Interestingly, the differences at are very small when keeping seeding and AGN feedback the same. The BHARD at cosmic noon versus at can vary somewhat, depending on accretion rate estimate, with a less pronounced peak and less dimming toward for less efficient accretion (in the free fall-based model in our case).
@@
The quasar luminosity functions are remarkably similar for the different accretion models. The main effect, most obvious in the occupation fractions at , is a non-linear interaction with AGN feedback. The transition to efficient AGN feedback, which differs for the different accretion models, has a more substantial effect on luminous AGNs than the inflow rate estimate (within the bounds tested in this work). In future works, a recalibration of the AGN feedback will be needed to match galaxy population constraints to properly assess the uncertainties in the MBH population. Notably, even prior to AGN feedback becoming a relevant factor for the galactic star formation rate, our results indicate an AGN feedback induced self-regulation of the accretion rate.
@@
The BH mass-stellar mass and BH mass-velocity dispersion relations change moderately, with a slight flattening of both relations for the free-fall based model. Notably, provided the overall normalisation is calibrated to the mass density, a weaker mass dependence of the accretion rate does not lead to substantially over-massive BHs at the low-mass end of the BH mass-stellar mass relation. This is a trend suggested by observations at high redshift (Pacucci et al. 2023). We speculate that a variable radiative or feedback efficiency would be required to achieve this, while making sure not to violate the constraints on the X-ray background.
We therefore conclude that the precise accretion rate formula used in galaxy formation simulations only has minor impacts on the galaxy and the SMBH populations once it is tuned to match the BHMD at . However, when we apply the different accretion rate estimates to a high-redshift simulation modelling an IMBH population with a seed mass around , drastic changes are evident. In particular, we find:
@@
Accretion rate densities differ by four orders of magnitude, while the mass growth is less affected due to growth via seeding. This indicates that our lack of understanding of the accretion mechanism leads to substantial uncertainties for possible signatures and growth histories of high redshift IMBHs.
@@
AGN luminosities differ by orders of magnitude depending on the model. We only show occupation fractions with luminosity thresholds below the detection limit of future X-ray surveys with Athena or AXIS (Marchesi et al. 2020). This is due to our relatively small simulation volume of Mpc side length that only contains low-mass IMBHs at this redshift. However, the trend indicates that slightly more massive BHs that would be present in larger volumes would be detectable at depending on accretion rate estimate. More works that include simulations of larger volumes are needed to explore whether these AGNs could constrain the inflow rate scaling onto MBHs. Such studies would also help us explore how sensitive these object are to variations in the quasar mode feedback.
These changes highlight the need to understand accretion, in particular, the BH mass dependence of the gas inflow to accurately model the high redshift IMBH regime. While we kept the seeding model fixed for the different SMBH simulations in this work, future simulations that model the cosmic evolution of MBHs from light seeds are also likely to show substantial accretion model dependent growth in SMBHs caused by the modified early growth history. Further studies combining this effort with effective seed models (e.g. Bhowmick et al. 2024b) are clearly needed to explore the possible IMBH evolution scenarios at high redshift and their consequences for luminous AGN and IMBH merger events, as well as the MBH population in the Local Universe. Further aspects missing from the employed model include: the treatment of (resolved) MBH dynamics, allowing for off-centre BHs, as well as recoil kicks from merging MBH binaries, which are expected to moderate the growth from accretion and mergers, respectively.
Overall, this new model represents a general gas inflow formula that parametrises different dependences on BH mass and different normalisations. It was developed to match the sprit of the ‘Learning the Universe’ project; specifically, it is meant to be applied as an effective model with parameters to be constrained by data or marginalised over in order to infer galaxy formation and cosmological parameters from structure formation. We deliberately chose to only use reliable gas properties in the simulation that are only weakly dependent on other model choices, such as in the employed ISM modelling (see Sivasankaran et al. 2022, for issues with the Bondi estimate). Measuring the simulation quantities at a fixed aperture also allows the model to be tested using high-resolution inflow simulations (Guo et al. 2023, 2024).